OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. What is (f ◦ g)(x) ? please describe the method you used to solve this problem.

jimthompson5910 (jim_thompson5910):

(f ◦ g)(x) means f(g(x))

jimthompson5910 (jim_thompson5910):

f(x) = 5x – 26 f(g(x)) = 5*( g(x) ) – 26 ... replace each x with g(x) f(g(x)) = 5*( 6x ) – 26 ... replace each g(x) on the right side with 6x now simplify

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw